OpenStudy (anonymous):

Please help... A photograph is 4 in. longer than it is wide. If the frame is 2 in. wide, the combined area of the photograph and the frame is 252 in^2. Find the dimensions of the photograph without the frame.

OpenStudy (anonymous):

Lowering the temperature _____ most chemical reactions. (1 point) slows down speeds up does not affect stops

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

Let Length be x width x-4 area of photgraph is x^2-4x

OpenStudy (anonymous):

how much bigger is the frame compared to the picture?

OpenStudy (anonymous):

It doesn't say

OpenStudy (anonymous):

X^2-4x<252 x^2-4x-252<0

OpenStudy (anonymous):

252 = (L+2)(W+2) = (L+2)(L-4 +2) = L² - 4 256 = L² L = 16 W = 12